Feb. 4-7

– Four stores provided ~37.25 hours of luminosity
– EE support replaced vacuum circuit breaker relays at Tevatron sectors D3 and A3
– NuMI vacuum circuit breaker repaired
– Further TEV VCB inspections halted beam to NuMI and Pbar
– Meson Cryo Central suffered power outage
– CHL reported a cold box failure
– Operations aborted store 8478
– Booster kicker cable repaired
– MTest experiment T-978 completed run

* The integrated luminosity for the period from 1/31/11 to 2/7/11 was 43.64 inverse picobarns. NuMI reported receiving 5.14E18 protons on target during this same period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts